Consider the cost function of Bamburi Cement Ltd given by Total cost (TC) = 4q2(squared)+ 16. Find the variable cost, fixed cost, average cost, average variable cost, average
fixed cost and the marginal cost for Bamburi Cement Ltd. ( 10 marks)
1
Expert's answer
2020-10-15T03:21:29-0400

We are given "TC = 4q^2 + 16"


Variable Cost (VC):

Variable costs are costs that vary with the level of output. Hence, they are deduced from the TC function by isolating the component affected by output (q).

"\therefore \space VC = 4q^2"


Fixed Costs (FC):

Fixed costs are costs not affected by level of production. They are constant at all production levels. Fixed costs are, thus, the independent component of the total cost.

"\therefore \space FC = 16"


Average Cost (AC):

"AC = \dfrac {TC}{q}"


"= \dfrac {4q^2 + 16}{q}"


"= 4q +\dfrac {16}{q}"


Average Variable Cost (AVC)

"AVC = \dfrac {VC}{q}"


"= \dfrac {4q^2}{q}"


"= 4q"



Average Fixed Cost (AFC):

"AFC = \dfrac {FC}{q}"


"= \dfrac {16}{q}"



Marginal Cost (MC):

"MC = \dfrac {d}{dq}(TC)"


"= \dfrac {d}{dq}(4q^2 + 16)"


"= 8q"